# A faster algorithm for Vertex Cover parameterized by solution size

### v6

@@ 1 Introduction
For an undirected graph $G=(V,E)$, a subset $S\subseteq V$ is called a vertex cover if every edge has at least one endpoint in $S$. It is closely related to an independent set, since if $S$ is an inclusion-wise minimal vertex cover then $V-S$ is an inclusion-wise maximal independent set, and vice-versa. Finding the size of the smallest vertex cover is a classic NP-complete problem [GJ79]. In particular, $G$ has a vertex cover of size at most $k$ if and only if it has an independent set of size at least $n-k$.
@@
subject to $\theta(u)+\theta(v)\geq 1$ for all edges $e=(u,v)\in E$
@@
The optimal solution to $\operatorname{LPVC}(G)$, denoted $\lambda(G)$ or just $\lambda$ if $G$ is clear from context, is a lower bound on the size of a minimum vertex cover of $G$. We also define $\mu(G)=k-\lambda(G)$, i.e. the gap between solution size and LP lower bound. This linear program has remarkable properties which have been exploited for a variety of algorithms [NT74, NT75, ${}^{+}$14]. For instance, an optimum basic solution is half-integral and can be found efficiently by a network flow computation.
[…]
@@ 1.1 Outline of our results
The FPT algorithms for vertex cover, including our new algorithm, are built out of recursively branching on high-degree vertices or other structures, while tracking some “measure” of the graph. Our main new idea is to use a measure which is a piecewise-linear function of $k$ and $\mu$. To illustrate, consider branching on whether some degree-$r$ vertex $u$ is in the cover: we recurse on the subproblem $G_{1}=G-u$ where $k$ is reduced by one and on the subproblem $G_{0}=G-u-N(u)$ where $k$ is reduced by $r$. We must also show that $\mu$ is significantly reduced in the two subproblems.
Suppose that, wishfully speaking, $\vec{\tfrac{1}{2}}$ remains the optimal solution to $\operatorname{LPVC}(G_{0})$ and $\operatorname{LPVC}(G_{1})$. This is what should happen in a “generic” graph. Then $\operatorname{\lambda}(G_{0})=|V(G_{0})|/2=(n-r-1)/2$ and $\operatorname{\lambda}(G_{1})=|V(G_{1})|/2=(n-1)/2$. Since $k_{0}=k-r$ and $k_{1}=k-1$, this implies that $\mu$ is indeed significantly decreased:
@@
But suppose, on the other hand, that $\vec{\tfrac{1}{2}}$ is not the optimal solution to $\operatorname{LPVC}(G_{0})$ or $\operatorname{LPVC}(G_{1})$. For a concrete example, suppose the neighborhood of $x$ in $G$ is a subset of that of $u$, so $G_{0}$ has an isolated vertex $x$ and an optimal solution $\theta^{*}$ to $\operatorname{LPVC}(G_{0})$ would set $\theta^{*}(x)=0$. In this situation, we develop an alternate branching rule for $G$: rather than branching on $u$ itself, we branch on the two subproblems where (i) both $u,x$ are in the cover and (ii) $x$ is not in the cover. See Figure 1 for an illustration.
Figure 1: In the subgraph $G-N[u]$, the vertex $x$ becomes isolated
We emphasize that this is just one example of our branching analysis. We need to handle more general situations where $\vec{\tfrac{1}{2}}$ is not the optimal solution to $\operatorname{LPVC}(G_{0})$ or $\operatorname{LPVC}(G_{1})$, which can require more complex branching rules. We emphasize that these branching rules are more powerful than simply using conventional branching rules along with the preprocessing rules; the latter may not be able to ensure a good reduction in $\mu$.
In Section 6, we use these ideas for a relatively simple algorithm with runtime $O^{*}(1.2575^{k})$, along with algorithms for bounded-degree graphs. This is already much better than [CKX10]. To improve further, and achieve Theorem 1, we need an additional idea: instead of branching on an arbitrary high-degree vertex, we branch on a targeted vertex to create additional simplifications in the graph. This is much more complex. For example, when we are branching on degree-4 vertices, it is favorable to choose one that has a degree-$3$ neighbor. This will create a degree-$2$ vertex in subproblem $\langle G-u,k-1\rangle$, leading to further simplifications of the graph. See Figure 2.
Figure 2: In the subgraph $G-u$, the vertex $v$ has degree two.
1.2 Related algorithms
Many different kinds of algorithms have been developed for the Vertex Cover problem. We cannot summarize this fully here; we describe some of the most relevant works for our own paper.
As we have mentioned, there is a long line of research into FPT algorithms parametrized by the solution size $k$. Currently, the fastest such algorithm has $O^{*}(1.2738^{k})$ runtime [CKX10]. In addition, [Tsu23] describes algorithms in graphs of maximum degree $3$ and $4$ with runtime respectively $O^{*}(1.1558^{k})$ and $O^{*}(1.2403^{k})$ respectively.
An important variant is Above-Guarantee Vertex Cover (AGVC), where the parameter is the difference between $k$ and various lower bounds on vertex cover size [MR99, MRS09, ${}^{+}$14, GP16, KKK22]. Of these, the parameter $\mu(G)=k-\operatorname{\lambda}(G)$ plays a particularly important role in this paper. We quote the following main result:
Theorem 2 ([${}^{+}$14]).
Vertex cover can be solved in time $O^{*}(2.3146^{\mu})$.
Theorem 3.
MaxIS-4 can be solved with runtime $O^{*}(1.137595^{n})$ by [XN17a].
MaxIS-5 can be solved with runtime $O^{*}(1.17366^{n})$ by [XN16]. [^2] This runtime is not claimed directly in [XN16], see [XN17b].
MaxIS-6 can be solved with runtime $O^{*}(1.18922^{n})$ by [XN17b].
MaxIS-7 can be solved with runtime $O^{*}(1.19698^{n})$ by [XN17b].
@@ 1.3 Notation
For a subset $S$ of $V$, the subgraphs of $G$ induced by $S$ and $V\setminus S$ are denoted by $G[S]$ and $G-S$, respectively. We write $u\sim v$ if $(u,v)\in E$ and $u\nsim v$ otherwise. For vertex sets $X,Y$, we write $X\sim Y$ if there exists $x\in X,y\in Y$ with $x\sim y$.
For a vertex $u$, the neighborhood $N_{G}(u)$ is the set $\{u\in V(G):u\sim v\}$ and the closed neighborhood $N_{G}[u]$ is the set $N_{G}(u)\cup\{u\}$. Extending this notation, for a vertex set $S\subseteq V(G)$, we write $N_{G}(S)=\bigl{(}\bigcup_{v\in S}N_{G}(v)\bigr{)}\setminus S$ and $N_{G}[S]=N_{G}(S)\cup S=\bigcup_{v\in S}N_{G}[v]$. For readability, we sometimes write $N(x,y)$ or $N[x,y]$ as shorthand for $N(\{x,y\})$ or $N[\{x,y\}]$.
The degree of vertex $v$, denoted by $\deg_{G}(v)$, is the size of $N_{G}(v)$. We call a vertex of degree $i$ an $i$-vertex; for a vertex $v$, we refer to a neighbor $u$ which has degree $j$ as a $j$-neighbor of $v$. An isolated vertex is a 0-vertex. We say a vertex is subquartic if it has degree in $\{1,2,3,4\}$ and subcubic if it has degree in $\{1,2,3\}$, and subquadratic if it has degree in $\{1,2\}$. (We do not count isolated vertices).
@@ 2 Preliminaries
We review some basic facts about branching and preprocessing rules for vertex cover. Much of this material is standard, see e.g. [${}^{+}$14]. We include proofs in Appendix A for completeness.
Our algorithm will heavily use the LP and its properties. This LP is closely related to properties of independent sets. For brevity, we refer to these as indsets, i.e. sets $X\subseteq V$ where no two vertices in $X$ are adjacent. For an indset $I$ in $G$, we define the surplus by $\operatorname{surp}_{G}(I)=|N_{G}(I)|-|I|$. If $G$ is clear from context, we write just $\operatorname{surp}(I)$ or $N(I)$. We define $\operatorname{minsurp}(G)$ to be the minimum value of $\operatorname{surp}_{G}(I)$ over all non-empty indsets $I$ in $G$, and a min-set to be any non-empty indset $I$ achieving this minimum, i.e. $\operatorname{surp}_{G}(I)=\operatorname{minsurp}(G)$.
Since this comes up in a number of places, we define $\operatorname{minsurp}^{-}(G)=\min\{0,\operatorname{minsurp}(G)\}$.
@@
$\lambda(G)=\tfrac{1}{2}(|V|+\operatorname{minsurp}^{-}(G)).$ Moreover, in a basic solution to $\operatorname{LPVC}(G)$, the set $I$ of vertices with value zero forms an indset (possibly empty) with $\operatorname{surp}_{G}(I)=\operatorname{minsurp}^{-}(G)$.
@@
For any critical-set $I$, there is a good cover $C$ with either $I\subseteq C$ or $I\cap C=\emptyset$.
@@
Proposition 6.
For any non-empty indset $I$ of $G$, there holds $\operatorname{minsurp}(G)\leq\operatorname{minsurp}^{-}(G-N[I])+\operatorname{surp}_{G}(I)$. The equality $\operatorname{minsurp}(G)=\operatorname{minsurp}^{-}(G-N[I])+\operatorname{surp}_{G}(I)$ holds if and only if $I$ is contained in a min-set of $G$.
Proposition 7.
The value $\operatorname{minsurp}(G)$ can be determined in polynomial time. Moreover, for any indset $X$ (possibly $X=\emptyset$), we can efficiently find a min-set of $G$ containing $X$ (if any such exists).
2.1 Preprocessing and branching rules
Our algorithm uses three major preprocessing rules. For this, we need to define the following graph structure: a funnel is a vertex $u$ with a neighbor $x$ such that $G[N(u)\setminus\{x\}]$ forms a clique. Here, we call $x$ the out-neighbor of $u$. For example, if a degree-3 vertex $u$ has a triangle with neighbors $t_{1},t_{2}$, and one other vertex $x$, then $u$ is a funnel with out-neighbor $x$; since this comes up so often, we refer to this as a 3-triangle. A well-known result is that, for a funnel $u$ with out-neighbor $x$, there exists a good cover $C$ with either $u\notin C$ or $x\notin C$.
@@ 2.1 Preprocessing and branching rules
Preprocessing Rule 3 (P3): Given a funnel $u$ with out-neighbor $x$, form graph $G^{\prime}$ with $k^{\prime}=k-1-\operatorname{codeg}(u,x)$ by removing vertices $N[u]\cap N[x]$ and adding edges between all vertices in $N(u)\setminus N[x]$ to all vertices in $N(x)\setminus N[u]$.
@@
These rules can be applied to any eligible sets in any order. Note that if a vertex $v$ has degree zero or one, then applying (P1) to $\{v\}$ removes $v$ and its neighbor (if any). If a vertex $v$ has degree two, then applying (P2) to $\{v\}$ removes $v$ and contracts its neighbors $x,y$ into a new vertex $v^{\prime}$. When no further preprocessing rules can be applied, we say the graph $G$ is simplified; the resulting graph has a number of nice properties, for instance, it has $\operatorname{minsurp}(G)\geq 2$, and the solution $\vec{\tfrac{1}{2}}$ is the unique optimal solution to $\operatorname{LPVC}(G)$, and the minimum degree of any vertex is $3$.
@@
There is one situation where rule P3 is particularly powerful. Consider a 3-vertex $u$ which has neighbors $x,y,z$ where $x\sim y$ and $y\sim z$. We refer to this as a kite; see Figure 4.
@@
Suppose $\operatorname{minsurp}(G)\geq 1$. Then applying (P3) to a kite in $G$ yields a graph $G^{\prime}$ with $k^{\prime}\leq k-2$ and $\mu(G^{\prime})\leq\mu(G)-1/2$.
2.2 Runtime and branching framework
Our algorithm follows the measure-and-conquer approach. Given an input graph $G$, it runs some preprocessing steps and then generates subproblems $G_{1},\dots,G_{t}$ such that $G$ is feasible if and only if at least one $G_{i}$ is feasible. It then runs recursively on each $G_{i}$. Such an algorithm has runtime $O^{*}(e^{\phi(G)})$ as long as the preprocessing steps have runtime $O^{*}(e^{\phi(G)})$ and it satisfies
$\sum_{i=1}^{t}e^{\phi(G_{i})-\phi(G)}\leq 1.$ (1)
(Here and throughout $e=2.718...$ is the base of the natural logarithm.) We refer to $\phi$ as the measure of the algorithm. For the most part, we will use measures of the form

### v7

@@ 1 Introduction
For an undirected graph $G=(V,E)$, a subset $C\subseteq V$ is called a vertex cover if every edge has at least one endpoint in $C$. It is closely related to an independent set, since if $C$ is an inclusion-wise minimal vertex cover then $V\setminus C$ is an inclusion-wise maximal independent set, and vice-versa. Finding the size of the smallest vertex cover is a classic NP-complete problem [GJ79]. In particular, $G$ has a vertex cover of size at most $k$ if and only if it has an independent set of size at least nkn-k.
@@
subject to $\theta(u)+\theta(v)\geq 1$ for all edges $(u,v)\in E$
@@
The optimal solution to $\operatorname{LPVC}(G)$, denoted $\lambda(G)$ or just $\lambda$ if $G$ is clear from context, is a lower bound on the size of a minimum vertex cover of $G$. We also define
$\mu(G)=k-\lambda(G),$
i.e. the gap between solution size and LP lower bound. This linear program has remarkable algorithmic properties [NT74, NT75, LNR+14]. For instance, an optimum basic solution is half-integral and can be found efficiently by a network flow computation.
Fixed-parameter tractable (FPT) and exact algorithms explore a landscape of parameters to understand the complexity of different problems [Nie06, FG06, DF99, CFK+15, FK10]. Vertex Cover was one of the first studied FPT problems with respect to the parameter $k$, the optimal solution size. Building on a long line of research, this culminated in an algorithm with $O^{*}(1.2738^{k})$ runtime [CKX10]. [^2] Throughout, we write $O^{*}(T)$ as shorthand for $T\cdot{\mathsf{poly}}(n)$. This record has been standing for more than a decade. Assuming the Exponential Time Hypothesis, no algorithm with runtime $O^{*}(2^{o(k)})$ is possible [CFK+15].
Many different kinds of algorithms have been developed for the Vertex Cover problem. We cannot summarize them fully here; we describe some of the most relevant works for our own paper.
An important variant is Above-Guarantee Vertex Cover (AGVC), where the parameter is the difference between $k$ and various lower bounds on vertex cover size [MR99, MRS09, LNR+14, GP16, KKK22]. Of these, the parameter $\mu$ plays a particularly important role in this paper. We quote the following main result:
Theorem 1 ([LNR+14]).
Vertex cover can be solved in time $O^{*}(2.3146^{\mu(G)})$.
Theorem 2.
MaxIS-4 can be solved with runtime $O^{*}(1.137595^{n})$ by [XN17a].
MaxIS-5 can be solved with runtime $O^{*}(1.17366^{n})$ by [XN16]. [^4] This runtime is not claimed directly in [XN16], see [XN17b] instead.
MaxIS-6 can be solved with runtime $O^{*}(1.18922^{n})$ by [XN17b].
MaxIS-7 can be solved with runtime $O^{*}(1.19698^{n})$ by [XN17b].
In addition, [Tsu23] describes algorithms in graphs of maximum degree $3$ and $4$ with runtime respectively $O^{*}(1.1558^{k})$ and $O^{*}(1.2403^{k})$ respectively.
@@ 1.1 Outline of our results
The FPT algorithms for vertex cover, including our new algorithm, are built out of recursive branching while tracking some “measure” of the graph. Our main new idea is to use a measure which is a piecewise-linear function of $k$ and $\mu$. To illustrate, consider branching on whether some degree-$r$ vertex $u$ is in the cover: we recurse on the subproblem $G_{1}=G-u$ with a solution of size $k_{1}=k-1$ and on the subproblem $G_{0}=G-u-N(u)$ with the solution size $k_{0}=k-r$. We must also show that $\mu$ is significantly reduced in the two subproblems.
Suppose that, wishfully speaking, $\theta=\vec{\tfrac{1}{2}}$ remains the optimal solution to $\operatorname{LPVC}(G_{0})$ and $\operatorname{LPVC}(G_{1})$. This is what should happen in a “generic” graph. Then $\operatorname{\lambda}(G_{0})=\frac{|V(G_{0})|}{2}=\frac{n-r-1}{2}$ and $\operatorname{\lambda}(G_{1})=\frac{|V(G_{1})|}{2}=\frac{n-1}{2}$. This implies that $\mu$ is indeed significantly decreased:
@@
But suppose, on the other hand, that $\vec{\tfrac{1}{2}}$ is not the optimal solution to $\operatorname{LPVC}(G_{0})$ or $\operatorname{LPVC}(G_{1})$. For a concrete example, suppose the neighborhood of $x$ in $G$ is a subset of that of $u$, so $G_{0}$ has an isolated vertex $x$ and an optimal solution $\theta^{*}$ to $\operatorname{LPVC}(G_{0})$ would set $\theta^{*}(x)=0$. In this situation, we develop an alternate branching rule for $G$: rather than branching on $u$ itself, we branch on the two subproblems where (i) both $u,x$ are in the cover and (ii) $x$ is not in the cover. (See Figure 1.)
Figure 1: In the subgraph $G_{0}$, the vertex $x$ becomes isolated.
We emphasize that this is just one example of our branching analysis. We need to handle more general situations where $\vec{\tfrac{1}{2}}$ is not the optimal solution to $\operatorname{LPVC}(G_{0})$ or $\operatorname{LPVC}(G_{1})$, which can require more complex and fundamentally new branching rules.
In Section 4, we use these ideas for a relatively simple algorithm with runtime $O^{*}(1.2575^{k})$, along with algorithms for bounded-degree graphs. This is already much better than [CKX10]. To improve further, and achieve Theorem 3, we need an additional idea: instead of branching on an arbitrary high-degree vertex, we branch on a targeted vertex to create additional simplifications in the graph. This is much more complex. For example, when we are branching on degree-4 vertices, it is favorable to choose one that has a degree-$3$ neighbor. This will create a degree-$2$ vertex in subproblem GuG-u, leading to further simplifications of the graph. See Figure 2.
Figure 2: In the subgraph GuG-u, the vertex $v$ has degree two.
@@ 1.2 Notation
For a subset $S$ of $V$, the subgraphs of $G$ induced by $S$ and $V\setminus S$ are denoted by $G[S]$ and GSG-S, respectively. We write $u\sim v$ if $(u,v)\in E$ and $u\nsim v$ otherwise. For vertex sets $X,Y$, we write $X\sim Y$ if there exists $x\in X,y\in Y$ with $x\sim y$.
For a vertex $u$, the neighborhood $N_{G}(u)$ is the set $\{v\in V(G):u\sim v\}$ and the closed neighborhood $N_{G}[u]$ is the set $N_{G}(u)\cup\{u\}$. Extending this notation, for a vertex set $S\subseteq V(G)$, we write $N_{G}(S)=\bigl(\bigcup_{v\in S}N_{G}(v)\bigr)\setminus S$ and $N_{G}[S]=N_{G}(S)\cup S=\bigcup_{v\in S}N_{G}[v]$. For readability, we sometimes write $N(x,y)$ or $N[x,y]$ as shorthand for $N(\{x,y\})$ or $N[\{x,y\}]$.
The degree of vertex $v$, denoted by $\deg_{G}(v)$, is the size of $N_{G}(v)$. We call a vertex of degree $i$ an $i$-vertex; for a vertex $v$, we refer to a neighbor $u$ which has degree $j$ as a $j$-neighbor of $v$. An isolated vertex is a 0-vertex. We say a vertex is subquartic if it has degree in $\{1,2,3,4\}$ and subcubic if it has degree in $\{1,2,3\}$, and subquadratic if it has degree in $\{1,2\}$.
@@
An independent set (or, for brevity, an indset) denotes a set $X\subseteq V$ where no two vertices in $X$ are adjacent, i.e. $X\not\sim X$.
@@ 2 Preliminaries
Our algorithm will heavily use the LP and its properties, which are closely related to properties of indsets. We review some basic facts here. Much of this material is standard, see e.g. [LNR+14]; we defer some proofs to Appendix A for completeness.
For an indset $I$ in $G$, we define the surplus by
$\operatorname{surp}_{G}(I)=|N_{G}(I)|-|I|.$
We define $\operatorname{minsurp}(G)$ to be the minimum value of $\operatorname{surp}_{G}(I)$ over non-empty indsets $I$ in $G$, and a min-set to be any non-empty indset $I$ with $\operatorname{surp}_{G}(I)=\operatorname{minsurp}(G)$. Since this plays a central role in our analysis, we define the shadow of a vertex set $X$, denoted $\operatorname{shad}_{G}(X)$, to be
$\operatorname{shad}_{G}(X)=\operatorname{minsurp}(G-X).$
We write $\operatorname{shad}_{G}(x_{1},\dots,x_{\ell},X_{1},\dots,X_{r})$ as shorthand for $\operatorname{shad}_{G}(\{x_{1},\dots,x_{\ell}\}\cup X_{1}\cup\dots\cup X_{r})=\operatorname{minsurp}(G-\{x_{1},\dots,x_{\ell}\}-(X_{1}\cup\dots\cup X_{r}))$. Since this comes up in a number of places, we define $\operatorname{minsurp}^{0}(G)=\min\{0,\operatorname{minsurp}(G)\}$. If $G$ is clear from context, we write just $\operatorname{surp}(I),N(I)$, etc.
To explain the notation, consider the following diagram:
Figure 3: In the subgraph GXG-X, there is an isolated vertex, hidden in the “shadow” of $X$. Here $\operatorname{shad}(X)=-1$.
@@
There holds $\lambda(G)=\tfrac{1}{2}(|V|+\operatorname{minsurp}^{0}(G)).$
Proposition 5.
Given any indset $I$ and vertex set $X\supseteq I$, one can efficiently find an indset which attains the minimum value of $\operatorname{surp}_{G}(J)$, over all indsets $J$ such that $I\subseteq J\subseteq X$.
Using Proposition 5, we can efficiently find min-sets of various types. For example, we can determine $\operatorname{minsurp}(G)$ by applying Proposition 5 with $X=V,I=\{v\}$ for all choices of vertex $v$; the minimum over all these values gives a (non-empty) min-set II of $G$. Similarly, we can find inclusion-wise minimal or maximal min-sets by repeatedly applying Proposition 5 to try to include or exclude each vertex from a given min-set. For the rest of the paper, we will not discuss the computational aspects of finding min-sets that are needed in our algorithm.
@@
For a critical-set $I$, there is a good cover $C$ with either $I\subseteq C$ or $I\cap C=\emptyset$. Moreover:
@@
2.1 Preprocessing rules and reductions in $k$
Our algorithm uses three major preprocessing rules. The first two are defined in terms of critical sets of small surplus.
@@ 2.1 Preprocessing rules and reductions in $k$
The final rule is based on a graph structure called a funnel. This structure consists of a vertex $u$ with a neighbor $x$ such that $G[N(u)\setminus\{x\}]$ forms a clique. We call $x$ the out-neighbor of $u$. For example, if a degree-3 vertex $u$ has a triangle with neighbors $t_{1},t_{2}$, and one other vertex $x$, then $u$ is a funnel with out-neighbor $x$; since this comes up so often, we refer to this as a 3-triangle. The following result is well-known:
Observation 7.
For a funnel $u$ with out-neighbor $x$, there exists a good cover $C$ satisfying either (i) $u\notin C$ or (ii) $x\notin C$ and $|C\setminus N(u)|=1$.
This leads to the following simplification rule:
Preprocessing Rule 3 (P3): Given a funnel $u$ with out-neighbor $x$, form graph $G^{\prime}$ with $k^{\prime}=k-1-\operatorname{codeg}(u,x)$ by removing vertices the vertices in $N[u]\cap N[x]$ and adding edges between all vertices in $N(u)\setminus N[x]$ to all vertices in $N(x)\setminus N[u]$.
@@
These rules can be applied to any eligible sets in any order. Note that if a vertex $v$ has degree zero or one, then applying (P1) to $\{v\}$ removes $v$ and its neighbor (if any). If a vertex $v$ has degree two, then applying (P2) to $\{v\}$ removes $v$ and contracts its neighbors $x,y$ into a new vertex $v^{\prime}$. When no further preprocessing rules can be applied, we say $G$ is simplified; the graph then has a number of nice properties, for instance, $\operatorname{minsurp}(G)\geq 2,\operatorname{mindeg}(G)\geq 3$, and the solution $\vec{\tfrac{1}{2}}$ is the unique optimal solution to $\operatorname{LPVC}(G)$.
@@
Observation 9.
Suppose $\operatorname{minsurp}(G)\geq 2$. Then any vertex set $X$ satisfies $\operatorname{shad}(X)\geq 2-|X|$, and any indset $I$ satisfies $\operatorname{shad}(N[I])\geq 2-\operatorname{surp}_{G}(I).$
In particular, a vertex $u$ satisfies $\operatorname{shad}(u)\geq 1$ and $\operatorname{shad}(N[u])\geq 3-\deg(u)$.
We define $S(G)$ to be the largest decrease in $k$ obtainable from applying rules (P1) – (P3) in some efficiently-computable sequence. [^5] It is not clear how to calculate the largest decrease in $k$ via preprocessing rules, since applying some rules may prevent other rules. We assume that we have fixed some polynomial-time computable sequence of potential preprocessing rules to reduce $k$ as much as possible. At various points in our algorithm, we will describe simplifications available for intermediate graphs. We always assume that our preprocessing rules can find such simplifications. For vertices $x_{1},\dots,x_{\ell}$ and vertex sets $X_{1},\dots,X_{r}$, we write $S(x_{1},\dots,x_{\ell},X_{1},\dots,X_{r})$ as shorthand for $S(G-\{x_{1},\dots,x_{\ell}\}-X_{1}-\dots-X_{r})$.
Proposition 10.
Let $x,y$ be non-adjacent subquadratic vertices. Then $S(G)\geq 2$ if any of the three conditions hold: (i) $\operatorname{codeg}(x,y)=0$ or (ii) $\deg(x)+\deg(y)\geq 3$ or (iii) $\operatorname{minsurp}(G)\geq 0$.
Proposition 11.
If $G$ has 2-vertices $x_{1},\dots,x_{\ell}$ with pairwise distance at least $3$, then $S(G)\geq\ell$.
Since a simplified graph $G$ has $\operatorname{minsurp}(G)\geq 2$, we have the following immediate observation from Proposition 10:
Observation 12.
If $G$ is simplified and $G-\{u,v\}$ has two non-adjacent subquadratic vertices, then $S(u,v)\geq 2$.
We now consider a more involved simplification rules.
Lemma 13.
If $\operatorname{minsurp}(G)\geq 0$ and we have an indset $I$ with $\operatorname{surp}_{G}(I)\leq 1$, then $S(G)\geq|I|$.
Proof.
Let $r=|I|$. If $\operatorname{surp}_{G}(I)=0$, or $I$ is a critical-set, we can apply (P1) or (P2) to II. So suppose that $\operatorname{surp}_{G}(I)=1$ and $\operatorname{surp}_{G}(K)=0$ for a non-empty subset $K\subsetneq I$. Choose $K$ to be inclusion-wise-maximal with this property, and let $s=|K|$. We apply (P1) to $K$, reducing $k$ by $s$ and obtaining a graph $G^{\prime}=G-N[K]$. Now consider the indset $J=I\setminus K$ in $G^{\prime}$, where |J|=rs|J|=r-s. We have $|N_{G^{\prime}}(J)|=|N_{G}(I)\setminus N_{G}(K)|=r+1-s$ and so $\operatorname{surp}_{G^{\prime}}(J)=1$.
We claim that $J$ is a critical-set in $G^{\prime}$. For, if some non-empty subset $K^{\prime}\subseteq J$ has $|N_{G^{\prime}}(K^{\prime})|\leq|K^{\prime}|$, then indset $I^{\prime}=K\cup K^{\prime}\subseteq I$ would have $|N_{G}(I^{\prime})|\leq|N_{G}(K)|+|N_{G^{\prime}}(K^{\prime})|\leq|K|+|K^{\prime}|=|I^{\prime}|$. Hence $\operatorname{surp}_{G}(I^{\prime})\leq 0$, contradicting maximality of $K$. So, we can apply (P2) to $J$ in $G^{\prime}$, getting a further drop of rsr-s. Overall, we get a net drop of s+(rs)=rs+(r-s)=r as desired. ∎
Lemma 14.
We define a kite to be a 3-vertex $u$ with neighbors $x,y,z$ where $x\sim y$ and $y\sim z$.
If $\operatorname{minsurp}(G)\geq 1$, then applying (P3) to a kite $u$ yields a graph $G^{\prime}$ with $k^{\prime}\leq k-2$ and $\mu(G^{\prime})\leq\mu(G)-1/2$. (See Figure 5.)
@@
Proof.
We view the application of (P3) to $u$ as a two-part process. First, we remove the shared neighbor $y\in N(u)\cap N(x)$; then, we merge $z,x$ into a new vertex $u^{\prime}$. The first step gives graph $G^{\prime\prime}=G-y$ with $k^{\prime\prime}=k-1,n^{\prime\prime}=n-1$. So, by Proposition 4, we have $\lambda(G^{\prime\prime})=\lambda(G)-1/2$ and so $\mu(G^{\prime\prime})\leq\mu(G)-1/2$. By Proposition 8, the second step gives $\mu(G^{\prime})\leq\mu(G^{\prime\prime})$ and $k^{\prime}=k^{\prime\prime}-1$. ∎
@@
Proof.
First, exhaustively apply (P1) to $G$; the resulting graph $G^{\prime}$ has $n^{\prime}\leq n,k^{\prime}\leq k$, and $\operatorname{minsurp}(G^{\prime})\geq 1$. In particular, $\mu(G^{\prime})=k^{\prime}-\lambda(G^{\prime})=k^{\prime}-n^{\prime}/2$. Note that $G^{\prime}\in\mathcal{G}$ since (P1) just deletes vertices. Now dovetail the two algorithms for $G^{\prime}$, running both simultaneously and returning the output of whichever terminates first. This has runtime $O^{*}(\min\{e^{cn^{\prime}},e^{a(k^{\prime}-n^{\prime}/2)+bk^{\prime}}\})$. For fixed $k^{\prime}$, this is maximized at $n^{\prime}=\frac{2k^{\prime}(a+b)}{a+2c}$, at which point it takes on value $e^{dk^{\prime}}\leq e^{dk}$. ∎
2.2 Branching framework
Our algorithm follows the measure-and-conquer approach. Given an input graph $G$, it runs some preprocessing steps and then generates subproblems $G_{1},\dots,G_{t}$ such that $G$ is feasible if and only if at least one $G_{i}$ is feasible. It then runs recursively on each $G_{i}$. Such an algorithm has runtime $O^{*}(e^{\phi(G)})$ as long as the preprocessing steps have runtime $O^{*}(e^{\phi(G)})$ and it satisfies [^6] Here and throughout $e=2.718$... is the base of the natural logarithm.
$\sum_{i=1}^{t}e^{\phi(G_{i})-\phi(G)}\leq 1$ (Branching Inequality)
We refer to $\phi$ as the measure of the algorithm. For the most part, we will use measures of the form